Caitlin is a co-op student. She invested her work-term earnings of $10,000 at 4% simple interest, paid annually. She intends to use the money in a few years to take a dream vacation with a friend.
(a) How long will the investment take to grow to a future value of $12,000?
(b) Determine Caitlin’s rate of return.

Answer by stanbon(75887)

Caitlin is a co-op student. She invested her work-term earnings of $10,000 at 4% simple interest, paid annually. She intends to use the money in a few years to take a dream vacation with a friend.
(a) How long will the investment take to grow to a future value of $12,000?
A(t) = P(1+rt)
12000 = 10,000(1+0.04t)
2000 = 10000*0.04t
0.2 = 0.04t
t = 5 years
------------
(b) Determine Caitlin’s rate of return.
Interest = 0.04*5*10000
I = 2000
Rate = 2000/5 = $400 per year
-------------------
Cheers,
Stan H.
-------